# An Efficient Learning Control Framework with Sim-to-Real for String-Type Artificial Muscle-Driven Robotic Systems

### v3

@@ VII-B Learning Control of Robotic Eye
Figure 9: Experimental trajectories of the foveation test of the robotic eye. Target angles are set at $(\theta^{*}=-5^{\circ},\psi^{*}=7.5^{\circ})$ and $(\theta^{*}=10^{\circ},\psi^{*}=-5^{\circ})$, respectively. (a), (c) The system output angles $(\theta,\psi)$. (b), (d) The corresponding actions $a_{t}$ generated by the trained actor.
Figure 10: Experimental results on the steady-state tracking errors in robotic eye at eyery combination of $(\theta^{*},\psi^{*})$ target angles spaced at $2.5^{\circ}$ intervals for $(|\theta^{*}|,|\psi^{*}|)\leq 10^{\circ}$ where the color bar represents the Euclidean distance norm of steady-state error.
@@
To demonstrate the applicability of the proposed learning control framework across diverse robotic systems driven by artificial muscle strings, this subsection presents experimental results on an additional platform, the 2-DOF robotic eye. We focus on the foveation control task in the experiment. The robotic eye, initialized at $\theta=\psi=0^{\circ}$ at ambient room temperature, is configured to foveate at various target angles $(\theta^{*},\psi^{*})$. Considering the system’s rise time, each testing episode lasts for $t_{\text{total}}=15$ s, with an action time step of $t_{a}=0.5$ s. The averaged Euclidean distance between the current and target angles within the last $5s$ window serves as the steady-state error metric, i.e., $e_{\text{ss}}=\sqrt{(\theta-\theta^{*})^{2}+(\psi-\psi^{*})^{2}}$.
Experimental trajectories with targets set at $(\theta^{*},\psi^{*})=(-5^{\circ},7.5^{\circ})$ and $(10^{\circ},-5^{\circ})$ are illustrated in Fig 9. These trajectories demonstrate that the measured angles rise to the target in approximately $5$ s, with negligible steady-state error. A similar field test is conducted with a series of target angles spaced at $2.5^{\circ}$ intervals within the defined range $|\theta^{*}|,|\psi^{*}|\leq 10^{\circ}$ to evaluate the overall control performance of the algorithm. The robotic eye is tested to foveate at each target in sequence. The results, illustrated in Fig. 10, reveal solid control performance, with steady-state errors consistently below $1^{\circ}$ for most target angles and an average of $0.56^{\circ}$.
In our prior research[20], we employed a DRL-based controller using the Deep Deterministic Policy Gradient algorithm[43], which was trained in simulation for 50,000 episodes. Additionally, a reduced-order state observer was used to estimate the real-time temperature of SCP muscles, achieving satisfactory control performance. In this paper, our proposed algorithm achieves comparable control performance with only 2,000 training episodes, representing just 1/25 of the original training time. Moreover, this is accomplished without the need to estimate real-time muscle temperature information under the condition of ambient temperature remaining relatively stable.
VII-C Randomization Ranges on Control Performance
@@
TABLE IV: Comparison of control errors between models trained under different randomization ranges
$\bar{e}_{\text{ss}}$ (deg) Randomization Range
Small Original Large
Real $0.63\pm 0.49$ $0.58\pm 0.37$ $0.76\pm 0.43$
Sim $0.38\pm 0.17$ $0.40\pm 0.17$ $0.74\pm 0.41$
In designing dynamics randomization, maintaining a balance between randomization level and model fidelity is crucial. Excessive randomization leads to an overly conservative controller, which may over-generalize and fail to leverage the more predictable aspects of system dynamics. Conversely, insufficient randomization causes the controller to overfit to a narrow set of conditions, thus limiting its performance when deployed in real-world systems. To investigate this trade-off, we conducted an experiment in which our control framework was trained with two additional randomization ranges—“small” and “large.” The small range was defined as half of the original randomization range, while the large range was doubled. We evaluated the performance of controllers trained under these varying randomization ranges in both simulated and real-world environments on the robotic wrist platform. The results, as summarized in Table IV, indicate that the control performance experiences a significant decline when randomization range is excessively large. Conversely, the smaller randomization range impairs the controller’s generalization capability from simulation to reality. Overall, a balanced range proves essential for ensuring robust performance.
@@ VIII Conclusion
This paper presents a learning-based control framework SAC-BAR designed for the string-type artificial muscle-driven robotic systems. To enhance training data efficiency and expedite convergence speed, we proposed two methods, namely bootstrap via PID and state-based data augmentation. Additionally, we incorporated randomization of artificial muscle dynamic parameters during training to address the sim-to-real gap for enhancing real-world control performance. Extensive experiments and ablation studies on two systems—a robotic eye driven and a parallel robotic wrist—validated the effectiveness and generalizability of the proposed control framework. In simulation, the designed bootstrap and augmentation methods reduced the required training episodes by approximately 60% compared to the baseline algorithm. Real-world experiments demonstrated that randomization of muscle dynamics improved the control performance of DRL-based controllers in targeted robotic systems, reducing the average error by 36.3%. Furthermore, our framework showed satisfactory performance on both robotic platforms, highlighting its potential applicability to similar artificial muscle robots.
Despite the strong performance of the proposed SAC-BAR, several limitations exist. The configuration of the reward function has a significant impact on experimental outcomes, often requiring empirical tuning to ensure optimal controller performance. The dynamic models of the artificial muscles and robotic wrist are simplified and do not fully represent their true dynamics. In future work, we plan to explore automated methods for reward function tuning, potentially incorporating adaptive or self-tuning mechanisms that respond dynamically to performance metrics. To better account for real-world variability, we will extend the randomization of muscle dynamics by incorporating high-fidelity models that capture nonlinearity and time-dependent factors, using data-driven or hybrid modeling approaches.

### v4

@@ VII-B Learning Control of Robotic Eye
Figure 10: Experimental trajectories of the foveation test of the robotic eye. Target angles are set at $(\theta^{*}=-5^{\circ},\psi^{*}=7.5^{\circ})$ and $(\theta^{*}=10^{\circ},\psi^{*}=-5^{\circ})$, respectively. (a), (c) The system output angles $(\theta,\psi,e_{\text{ss}})$. (b), (d) The corresponding actions $a_{t}$ generated by the trained actor.
@@
To demonstrate the applicability of the proposed learning control framework across diverse robotic systems driven by artificial muscle strings, this subsection presents experimental results on an additional platform. The robotic eye, initialized at $\theta=\psi=0^{\circ}$ at ambient room temperature, is configured to foveate at various target angles $(\theta^{*},\psi^{*})$. Considering the system’s rise time, each testing episode lasts for $t_{\text{total}}=15$ s, with an action time step of $t_{a}=0.5$ s. The averaged Euclidean distance between the current and target angles within the last $5s$ window serves as the steady-state error metric, i.e., $e_{\text{ss}}=\sqrt{(\theta-\theta^{*})^{2}+(\psi-\psi^{*})^{2}}$. Experimental trajectories with targets set at $(\theta^{*},\psi^{*})=(-5^{\circ},7.5^{\circ})$ and $(10^{\circ},-5^{\circ})$ are illustrated in Fig 10. These trajectories demonstrate that the measured angles rise to the target in approximately $5$ s, with negligible steady-state error. A similar field test is conducted to evaluate the overall control performance of the algorithm. The results, illustrated in Fig. 11, reveal solid control performance, with steady-state errors consistently below $1^{\circ}$ for most target angles and an average of $0.56^{\circ}$. Our prior research[17] employed a Deep Deterministic Policy Gradient-based controller[43], which was trained in simulation for $50,000$ episodes. In this paper, our proposed algorithm achieves comparable control performance with only $2,000$ training episodes, representing just $1/25$ of the original training time.
Figure 11: Experimental results on the steady-state errors in robotic eye at eyery combination of $(\theta^{*},\psi^{*})$ target angles spaced at $2.5^{\circ}$ intervals for $(|\theta^{*}|,|\psi^{*}|)\leq 10^{\circ}$.
VII-C Different Randomization Variances
@@
TABLE IV: Comparison of control errors for models trained with different variances of the randomization distribution
$\bar{e}_{\text{ss}}$ (deg) Standard Deviation Values
$0.5\sigma_{0}$ $1.0\sigma_{0}$ $1.5\sigma_{0}$ $2.0\sigma_{0}$
Sim 0.38±$\pm 0.17$ 0.40±$\pm 0.17$ 0.60±$\pm 0.35$ 0.74±$\pm 0.41$
Real 0.63±$\pm 0.49$ 0.58±$\pm 0.37$ 0.64±$\pm 0.38$ 0.76±$\pm 0.43$
To evaluate the influence of randomization variance on control performance, we conduct experiments using four standard deviation (SD) settings: $0.5\sigma_{0}$, $1.0\sigma_{0}$, $1.5\sigma_{0}$, and $2.0\sigma_{0}$, where $\sigma_{0}$ represents the baseline SD. Controllers trained under each randomization setting are evaluated in both simulated and real-world environments using the robotic wrist platform. As summarized in Table IV, the controller trained with a $2.0\sigma_{0}$ range exhibits a significant degradation in performance across both domains. In contrast, the controller trained with a $0.5\sigma_{0}$ range demonstrates limited generalization, failing to accommodate the variability inherent in real-world dynamics. While the selection of randomization variance significantly affects the resultant control performance, all controllers successfully converge during training. This indicates that the proposed method ensures reliable convergence across a reasonably bounded range of randomization settings. These results underscore the necessity of carefully tuning the randomization variance to balance generalization and performance without compromising convergence stability.
@@ VIII Conclusion
This paper presents a learning control framework SAC-BAR designed for the string-type artificial muscle-driven robotic systems. To enhance training data efficiency, we proposed two methods, namely bootstrap via PID and state-based data augmentation. Additionally, we incorporated randomization of artificial muscle dynamic parameters during training to address the sim-to-real gap for enhancing real-world control performance. Extensive experiments and ablation studies on two robotic systems—a robotic eye and a parallel robotic wrist—validated the effectiveness and generalizability of the proposed control framework. In simulation, the designed bootstrap and augmentation methods reduced the required training episodes by approximately 60% compared to the baseline algorithm. Real-world experiments demonstrated that randomization of muscle dynamics improved the control performance in targeted robotic systems, reducing the average error by 36.3%. Furthermore, our framework demonstrated satisfactory performance on both robotic platforms, highlighting its potential applicability to other robotic systems with actuators characterized by high nonlinearity and parameter uncertainties.
Despite the strong performance of the proposed SAC-BAR, several limitations exist. The configuration of the reward function relies on empirical tuning, and this paper does not provide a formal theoretical convergence analysis under muscle dynamic randomization. Additionally, the application scenarios explored in this work are relatively limited, and the robotic testing platforms utilized share similar characteristics. In future work, we aim to explore automated methods for reward function tuning and develop theoretical convergence analysis frameworks that consider bounded randomization ranges and non-stationary dynamics. To further evaluate the practicality of the proposed method, we plan to extend its application to more complex systems and diverse scenarios.